OpenStudy (anonymous):

Solve using substitution {2y=x+2 {6x-12y=0

OpenStudy (chihiroasleaf):

2y=x+2 x=2y-2 substitute x=2y-2 into 6x-12y=0, can you continue it? :)

OpenStudy (anonymous):

answer is no solution because in the end it becomes -12 =0 so yea

OpenStudy (anonymous):

so no solution is the final answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay thank you!